OpenStudy (anonymous):

can someone please help me..... write the slope-intercept equation for the line that passes through (-4,6)and (2,-6)..plz help me its due in a few hours

OpenStudy (phi):

first find the slope m= (y1-y2)/(x1-x2)

OpenStudy (mertsj):

Find the slope. m=(-6-6)/(2-(-4)) m=-12/6 m=-2

OpenStudy (mertsj):

Now choose either point and plug into point-slope form

OpenStudy (anonymous):

huh?

OpenStudy (mertsj):

I choose (-4,6) y-6=-2(x-(-4)) y-6=-2(x+4) y-6=-2x-8 y=-2x-2

OpenStudy (mertsj):

And that is slope-intercept form

OpenStudy (anonymous):

so that how i write it

OpenStudy (mertsj):

Point slope form: